OpenStudy (anonymous):

x/2 + y/8 = 4 x/3+y/2=-2 Elimination , please help! :)

OpenStudy (anonymous):

maybe clear the fractions first

OpenStudy (anonymous):

im not getting the right answer :(

OpenStudy (anonymous):

\[\frac{x}{2}+\frac{y}{8}=4\] \[4x+y=32\] for the first one

OpenStudy (anonymous):

\[x/3+y/2=-2\] \[2x+3y=-12\] for the second one then work with \[4x+y=32\] \[4x+3y=-12\]

OpenStudy (anonymous):

\[-4x-y=-32\] \[4x+3y=-12\] \[2y=-44\] \[y=-22\] does that look right ?

OpenStudy (radar):

When you multiplied the "second" one by two (to get the 4x) why did you not multiply all the terms?

OpenStudy (anonymous):

i think i multipied the second one by 6 at least that is what i meant to do

OpenStudy (anonymous):

oh i see i messed up didn't i?

OpenStudy (radar):

Ooh kay

OpenStudy (anonymous):

no no i made a mistake

OpenStudy (anonymous):

\[4x+y=32\] \[2x+3y=-12\]should be the system i messed up

OpenStudy (anonymous):

btw did you eliminate x or y?

OpenStudy (radar):

Which one do you want to eliminate?

OpenStudy (anonymous):

y

OpenStudy (anonymous):

do i multiply the first one by 1 and the second one by 3?

OpenStudy (radar):

You need to make the coefficients of y the same.

OpenStudy (anonymous):

i tried but it didnt divide evenly

OpenStudy (radar):

Yes, but no need to multiply the first one by one as it would simply remain the same.

OpenStudy (radar):

You are on the right track, gotta run.

OpenStudy (anonymous):

x-3y=10 12+3=36 13x = 16

OpenStudy (anonymous):

it dosent divide evenly

OpenStudy (anonymous):

13x=46**

OpenStudy (radar):

@kellyking, First and foremost you must give up the notion that correct answers are going to be "even" Sattelite73 showed you how those equations became equal to: 1. 4x + y = 32 and 2. 2x + 3y = -12 You elected to eliminate the y. To do this multiply equation 1 by -3, getting: 1. -12x - 3y =-96 now add equation 2 2. 2x +3y = -12 --------------------- -10x = -108 x= 108/10 = 54/5 Not a pretty answer .....but correct. Now substitute this value of x (54/5) into equation 2. Getting 2(54/5) + 3y = -12 108/5 + 3y = -12 multiply thru by 5 getting: 108 +15y = -60 15y=-60-108 15y=-168 y=-168/15=-56/5 another odd looking value, but none the less correct solution. Sorry I had to run last night, but Thurs is my favorite TV night lol.